Why power is directly as well inversely proportional to resistance.??

Grade:10

1 Answers

vinay gupta
6 years ago
We know that ,power is equal to multiplection of voltage and current (P=V*I).so current is equal to devidection of voltage and resistance (I=V/R). so power is (P=V^2/R).so power is inversely proportional to resistance.